### v2

# Personalized Image Generation for Recommendations Beyond Catalogs

Abstract. Personalization is central to human-AI interaction, yet current diffusion-based image generation systems remain largely insensitive to user diversity. Existing attempts to address this often rely on costly paired preference data or introduce latency through Large Language Models. In this work, we introduce REBECA (REcommendations BEyond CAtalogs), a lightweight and scalable framework for personalized image generation that learns directly from implicit feedback signals such as likes, ratings, and clicks. Instead of fine-tuning the underlying diffusion model, REBECA employs a two-stage process: training a conditional diffusion model to sample user- and rating-specific image embeddings, which are subsequently decoded into images using a pretrained diffusion backbone. This approach enables efficient, fine-tuning-free personalization across large user bases. We rigorously evaluate REBECA on real-world datasets, proposing a novel statistical personalization verifier and a permutation-based hypothesis test to assess preference alignment. Our results demonstrate that REBECA consistently produces high-fidelity images tailored to individual tastes, outperforming baselines while maintaining computational efficiency [^1] Code available in anonymized repository during review..
1 Introduction
Personalization is central to human-AI interaction, as users exhibit diverse tastes, intents, and creative goals. However, current diffusion-based image generation systems remain largely insensitive to such user diversity, producing outputs that are visually impressive but not tailored. Addressing this limitation is crucial for advancing applications in creative tools, advertising, and content recommendation.
Recent work has begun to address personalization in diffusion models. Personalized Preference Fine-tuning (PPD) [dang2025personalized] adapts diffusion models to user tastes through pairwise preference supervision, where users choose preferred images from pairs. While effective, this approach depends on costly paired data and model fine-tuning, limiting scalability. Personalized Multimodal Generation (PMG) [shen2024pmg] instead leverages large language models (LLMs) to infer user preferences from behaviors such as clicks or chat histories, but introduces latency from LLM inference and relies heavily on textual cues rather than richer multimodal signals.
In this work, we introduce REBECA (REcommendations BEyond CAtalogs), a lightweight and scalable framework for personalized image generation with diffusion models. Unlike prior methods that rely on paired annotations or large language models, REBECA learns personalization directly from behavioral data such as likes, ratings, and clicks that are readily available in real-world social media platforms. The method operates in two stages:
  1. 1.
    Train a conditional diffusion to model the distribution $p_{\hat{\theta}}(I_{e}\mid U,R)$. Then samples image embeddings $I_{e}$ conditioned on a user $U$ and rating $R$.
  2. 2.
    Decode $I_{e}$ into personalized images $I$ using a pretrained diffusion model $p(I\mid I_{e})$ that conditionally decodes the embedding.
Unlike prior work, REBECA does not rely on textual descriptions, paired preference labels, or per-user fine-tuning. Instead, we learn a single conditional diffusion prior whose geometry spans the preference manifolds of all users. Conditioning identifies a user-specific region of this shared space, enabling scalable, plug-and-play personalization
We rigorously evaluate REBECA on real-world user datasets, measuring both personalization strength and visual fidelity. In addition to more traditional metrics such as recall, precision, and predicted image quality, we propose and employ a statistical personalization verifier and a permutation hypothesis test for personalized generations. Our results on synthetic and real data show that REBECA consistently produces images that align with individual user preferences. In summary, our contributions are threefold:
  1. 1.
    Learning from implicit signals. We introduce a single user-conditioned diffusion prior that models personalized CLIP-space embeddings directly from implicit feedback, eliminating the need for preference pairs, LLM mediation, or per-user fine-tuning.
  2. 2.
    Lightweight, plug-and-play framework. REBECA decouples personalization from the image generator: the learned prior produces embeddings that plug into any pretrained decoder with the same embedding class, enabling large-scale personalized generation with minimal compute.
  3. 3.
    Rigorous evaluation. We propose a statistical personalization verifier and a permutation-based hypothesis test to assess alignment with user preferences, complementing standard metrics, and show strong performance on both synthetic and real datasets. We believe our newly introduced evaluation procedure can be used by future work as a standard evaluation approach for personalized generations.
2 Related Work
Figure 2: REBECA overview. Training: Conditional diffusion prior trained to generate personalized image embeddings from user IDs and ratings. Inference: Generated embeddings are decoded into images via a pretrained image decoder model.
Parameter-Efficient Fine-Tuning (PeFT). Adapters have emerged as an efficient means of specializing large text-to-image models without updating the full backbone [hu2021loralowrankadaptationlarge, zhang2023adding, ye2023ip-adapter]. In diffusion models, IP-Adapter [ye2023ip-adapter] extends controllability to visual references by injecting image-conditioned cross-attention layers, while methods such as LoRA, Textual Inversion, and DreamBooth [hu2021loralowrankadaptationlarge, gal2022image, ruiz2023dreambooth, mou-t2i] provide instance-level personalization from a handful of example images. However, these approaches require explicit reference content (captions or images) and therefore do not apply directly to implicit-feedback or recommendation settings where preferences must be inferred rather than provided as exemplars.
REBECA generalizes the adapter concept by conditioning the diffusion prior on user embeddings learned from behavioral data, rather than textual or visual exemplars. This yields a compact, reusable personalization layer that captures user-level preference structure and scales across hundreds of users without per-user fine-tuning.
Diffusion Model Alignment via Reinforcement Learning. A complementary line of work aligns diffusion models to external objectives using reinforcement learning. Several methods apply PPO-style optimization to diffusion trajectories or denoising steps [black2024trainingdiffusionmodelsreinforcement, fan2023dpokreinforcementlearningfinetuning, zhang2024largescalereinforcementlearningdiffusion], using rewards from aesthetic models, human preference predictors, or task-specific heuristics. Other approaches introduce reward-conditioned diffusion objectives [zhao2025addingconditionalcontroldiffusion], or train reward models that guide sampling in a policy-improvement loop.
These methods operate on text prompts or global image-level rewards and typically require dense, high-quality reward signals. In contrast, REBECA learns from implicit feedback and models personalized embedding distributions directly, providing user-level alignment without policy iteration or reinforcement learning.
Personalized Preference Alignment. More recent work introduces explicit preference alignment. Personalized Preference Fine-tuning of Diffusion Models (PPD) [dang2025personalized] extends Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage] to the multi-user setting. PPD derives user embeddings from the hidden representations of a vision–language model (VLM) trained on a few pairwise preference examples per user. It then fine-tunes a decoupled cross-attention module under multiple reward objectives to align generation with those embeddings. While effective, PPD relies on paired preference data, repeated VLM evaluations per user, and additional sample generation for each forward pass, which severely limits scalability.
Personalized Multimodal Generation (PMG) [shen2024pmg], leverages large language models (LLMs) to infer user preferences from behavioral traces such as images, clicks, or conversations. PMG translates these multimodal behaviors into natural-language prompts, keywords, and embedding representations that condition a separate diffusion-based generator. However, this approach introduces latency and computational cost due to the LLM inference stage, and personalization remains largely mediated by text rather than by directly learning aesthetic alignment.
REBECA differs fundamentally from these approaches: it dispenses with both preference-pair supervision and VLM-based mediation, learning personalization directly from implicit feedback signals such as likes, ratings, or clicks. By conditioning a diffusion prior on compact user embeddings, REBECA achieves scalable, plug-and-play personalization while preserving the generality of pretrained image generators.
3 Methodology
We present the REBECA pipeline, a two-stage framework for personalized image generation from implicit feedback. In training, images are first mapped to a lower-dimensional embedding space using CLIP [radford2021learningtransferablevisualmodels]. A conditional diffusion model [ho2020denoisingdiffusionprobabilisticmodels] is then trained on these embeddings using classifier-free guidance [ho2022classifierfreediffusionguidance], conditioned on user and rating information. At inference time, the trained embedding diffusion model is provided with a user ID and a desired rating to generate image embeddings, which are then decoded into images using a pre-trained image generator. REBECA consists of two key components:
  1. 1.
    Personalized embedding generator. A conditional diffusion prior samples CLIP-space embeddings directly conditioned on user IDs and ratings.
  2. 2.
    Decoder model. The generated embeddings are translated into images by a frozen text-to-image model (e.g., Stable Diffusion  [podell2023sdxlimprovinglatentdiffusion, rombach2021highresolution]) augmented with an IP-Adapter [ye2023ip-adapter] to enable image-conditioned generation.
See Figure 2 for an overview. These design choices address the limitations of prompt-based personalization and enable flexible, language-free conditioning in CLIP space.
Personalized image generation. We aim to sample personalized images $I$ from $p(I\mid U,R)$, where $U$ denotes a user and $R$ a desired rating. We factorize:
$p(I\mid U,R)$ $=p(I,I^{e}\mid U,R)$
$=p(I^{e}\mid U,R)\,p(I\mid I^{e},U,R),$ (1)
where $I^{e}$ is the deterministic CLIP embedding of $I$. We approximate the two conditional distributions as follow. First, a conditional diffusion prior $p_{\theta}(I^{e}\mid U,R)$ is trained to approximate $p(I^{e}\mid U,R)$ and generate personalized embeddings:
$I^{e}\mid U,R\sim p_{\hat{\theta}}(I^{e}\mid U,R).$ (2)
Second, a pretrained text-to-image model with an IP-Adapter $p(I\mid I^{e})$ approximates $p(I\mid I^{e},U,R)$ and decodes these embeddings into visual samples:
$I\mid I^{e}\sim p(I\mid I^{e}).$ (3)
This approximation assumes that CLIP embeddings retain the information relevant for user preference; prior work shows that CLIP spaces exhibit smooth semantic directions and well-structured manifolds [radford2021learningtransferablevisualmodels, ramesh2022hierarchicaltextconditionalimagegeneration, levi2025the], making them a suitable domain for modeling $p(I^{e}\mid U,R)$. Our key insight is that user preferences, when expressed as implicit feedback, naturally define regions of CLIP-space that can be learned as a conditional diffusion prior.
Conditional diffusion prior. Let $I^{e}\in\mathbb{R}^{d}$ denote a CLIP image embedding. We model the conditional prior $p_{\theta}(I^{e}\mid U,R)$ via a standard forward diffusion process:
$q(I_{t}^{e}\mid I_{0}^{e})$ $=\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{t}}\,I_{0}^{e},\,(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (4)
with $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The denoising network $f_{\theta}$ is trained to predict the clean embedding $I_{0}^{e}$ directly from a noisy sample:
$\mathcal{L}_{x_{0}}(\theta)=\mathbb{E}_{I_{0}^{e},\,U,R,\,t,\,\epsilon}\!\left[\big\|f_{\theta}(I_{t}^{e},t,U,R)-I_{0}^{e}\big\|_{2}^{2}\right],$ (5)
where $\epsilon\!\sim\!\mathcal{N}(0,\mathbf{I})$ and $t\!\sim\!\mathrm{Uniform}\{1,\ldots,T\}$. At inference, classifier-free guidance (CFG) is applied as
$f_{\theta}^{(s)}=f_{\theta}^{(0)}+\omega\!\left(f_{\theta}^{(1)}-f_{\theta}^{(0)}\right),$ (6)
where $f_{\theta}^{(1)}=f_{\theta}(I_{t}^{e},t,U,R)$ and $f_{\theta}^{(0)}=f_{\theta}(I_{t}^{e},t,\varnothing,\varnothing)$. The guidance scale $\omega$ controls personalization strength. The final denoised embedding $\hat{I}_{0}^{e}$ is decoded into an image using the frozen decoder, yielding personalized samples consistent with the target user and rating.
REBECA Prior Architecture. Our diffusion prior is intentionally lightweight: a 4.44.4M-parameter transformer with 6 PriorBlocks (Appendix A), AdaLN-Zero conditioning, and a learned tokenizer that splits CLIP embeddings into tokens. Conditioning enters through user, rating, and timestep tokens combined via a small MLP. The small scale allows end-to-end training in <10<10 minutes on a single RTX-4090, enabling rapid iteration and scalability to large user sets. A full description of our training protocol, and final configuration may be found in Appendix B.
4 REBECA Under Control
We design a controlled simulation to evaluate REBECA when user preferences are known.
4.1 Dataset and Setup
We construct a controlled setting using dSprites [dsprites17], restricting to red/blue hearts and squares (four shape–color pairs). Four synthetic users each prefer one pair $(s_{U},c_{U})$, and we generate ratings $R\in\{0,1\}$ via a simple probabilistic rule that assigns high probability ($0.95$) to the preferred pair, low probability ($0.05$) to the opposite, and intermediate probability ($0.10$) to mismatched shape or color. We sample 40,00040{,}000 rated images and split them $90/10$.
4.2 Image Generation
REBECA is implemented using a lightweight Variational Autoencoder (VAE) [VAE-2019, kingma2022autoencodingvariationalbayes] with a 32-dimensional latent space (Figure 3) trained on the dSprites dataset. We construct a small VAE solely to define a ground-truth latent space with a frozen decoder that reconstructs final images from embeddings.. The encoder provides compact representations of dSprites, and the learned prior generates new embeddings aligned with user taste.
Figure 3: UMAP [umapMcInnes2018] projection of VAE embeddings. Color and shape clusters are cleanly separated.
4.3 Evaluation
Evaluation proceeds in two stages. (1) Qualitative: UMAP projections visualize whether user embeddings cluster coherently according to preferences. (2) Quantitative: REBECA is compared with (a) a mean-embedding baseline that generates a single prototype image per user and (b) a random baseline sampling uniformly from training data. For each user, REBECA and the random baseline generate $k{\in}\{1,5,10,20,25\}$ samples. We compare REBECA to (a) a mean-embedding baseline that generates a single prototype per user, and (b) a random baseline that samples uniformly from the training set. For $k\in\{1,5,10,20\}$, we report Precision@k (fraction of generated images whose nearest neighbor in the test set belongs to the user’s liked subset) and Recall@k (fraction of liked test images for which at least one generated sample is the nearest neighbor).
Table 1: Precision@k and Recall@k averaged over users in the controlled simulation. Recall is scaled by $\times 10^{-3}$. The mean baseline is deterministic and thus lacks diversity.
Method $k$ Precision@$k$ Recall@$k$ ($\times 10^{-3}$)
REBECA 1 0.782 0.167
5 0.782 0.835
10 0.789 1.685
20 0.799 3.409
Mean (deterministic) 1 1.000 0.214
Random 1 0.290 0.062
5 0.282 0.301
10 0.287 0.613
20 0.291 1.244
User 0 User 11 User 22 User 33
Likes Refer to caption Refer to caption Refer to caption Refer to caption
Generations Refer to caption
Figure 4: Per-user visualization in the controlled setting. Top: liked samples for each user. Bottom: images generated by REBECA using the frozen VAE decoder. REBECA captures each user’s preference manifold while maintaining diversity.
4.4 Results
Figure 4 shows that generated samples align with users’ liked regions in latent space. Table 1 confirms that REBECA achieves the best precision-recall balance, capturing user tastes while preserving diversity.
This controlled setting validates that REBECA recovers user-specific preference manifolds when ground truth is known. In Section  Sec. 5 we show that the same structure appears in real-world behavioral data, using a learned verifier as a surrogate for unknown preferences.
5 REBECA In The Wild
In this section, we evaluate REBECA against strong baselines in a realistic scenario, using images rated by real users.
Figure 5: Comparison of personalization performance across generation approaches. REBECA achieves the highest user scores, surpassing VLM-based baselines and LoRA fine-tuning variants.
5.1 Dataset and Setup
We employ the FLICKR-AES dataset [Ren_2017_ICCV]. The dataset contains a curated sample of $40,988$ Creative Commons-licensed photos from FLICKR, rated for aesthetic quality on a scale of 1 to 5. Annotations were collected using Amazon Mechanical Turk, with each image receiving ratings from five different workers. In total, 210 workers contributed, scoring 40,988 images for a total of $193208$ ratings.
Thus, our data consist of triplets of users, ratings, and images $(U,R,I)$. CLIP embeddings $I^{e}$ are deterministic mappings of images, and we obtain them with a single pass through the appropriate CLIP ViT-H-14 [cherti2023reproducible] encoder. We further map the ratings to binary ratings, representing like and dislike: given a pair of user and a rated image, we let $R=1$ if the rating is $4$ or higher, and $R=0$ otherwise. At the training stage, we fit our image embedding prior with $(U,R,I^{e})$. Importantly, image captions are not part of our data, and we uniquely rely on users’ historical ratings.
5.2 Image Generation
To ensure that the sampled embeddings align with user preferences, we set $R=1$ to indicate a strong preference signal, thereby prioritizing images that align with previously liked content. At first sight, it might appear that training on liked and disliked images to generate only positive samples is unnecessary. However, this decision is principled, and the reasoning is twofold. First, learning from both positive and negative feedback enables the model to build more informative user embeddings, capturing a fuller picture of individual preferences. Second, there is strong empirical evidence that conditional generative models consistently outperform their unconditional counterparts in terms of sample quality and alignment [dhariwal2021diffusionmodelsbeatgans, brock2019largescalegantraining].
5.3 Evaluation
All image generation uses Stable Diffusion v1.5 [podell2023sdxlimprovinglatentdiffusion, rombach2021highresolution] as a frozen decoder, operating at $512{\times}512$ resolution in fp16. We generate $25$ images per user using $50$ denoising steps with classifier-free guidance (CFG) set to $5$, balancing personalization and diversity. We used empty system prompts during these experiments. The diffusion prior outputs a personalized embedding $I^{e}$, which is then provided to the SD 1.5 pipeline equipped with the IP-Adapter [ye2023ip-adapter].
5.3.1 Baselines
Baselines require image captions. To obtain semantic descriptions of the training images, we use the LLaVA-1.5-7B multimodal model [liu_llava] as an image-language tagger. Each image is passed through LLaVA with a fixed JSON-style instruction that requests a concise caption together with short lists of objects, attributes, styles, and colors. See Appendix C for a more detailed specification. The model outputs are parsed into a structured table of captions and visual keywords used downstream for preference modeling. Next, we introduce the baselines in more detail.
LoRA per User. We train a separate LoRA adapter [hu2021loralowrankadaptationlarge] for each user on that user’s Flickr subset, keeping the CLIP text encoder frozen. Hyperparameters (rank, steps, warmup) are adapted to the number of images per user (Sec. D.1).
Shared LoRA. A single LoRA adapter is trained on the union of all users, then loaded into a fresh SD1.5 pipeline at evaluation time. See Sec. D.2.
VLM Personas. Following PMG-style text conditioning [shen2024pmg], we derive per-user textual personas from LLaVA-generated tags, including short descriptions and positive/negative keyword lists. Prompts are built by combining persona text with sampled positive keywords and standard degradation terms in the negative prompt. We refer the reader to Sec. D.3 for a comprehensive description.
VLM Personas + LoRA. We also combine per-user LoRA with VLM personas, loading each user’s LoRA adapter and prompting with their persona during generation.
Refer to caption Refer to caption
(a) LAION aesthetic predictor (b) HPSv2 overall quality predictor
Figure 6: Quantitative comparison of aesthetic and quality predictors across personalization methods. (a) The LAION aesthetic predictor captures visual appeal independently of textual prompts. (b) The HPSv2 predictor measures prompt-conditioned human preference and image quality. Across both metrics, REBECA consistently achieves the best performance.
5.3.2 Metrics
In addition to precision@$k$ and recall@$k$ (Section 4.3), we introduce additional metrics for quantitative evaluation.
Personalization Verifier. Since we cannot directly evaluate the preferences of users from the FLICKR-AES dataset on generated images, we cannot immediately determine whether the images produced by REBECA align with individual tastes. To address this, we train a predictive model that estimates the probability that a particular user $U$ will like a given image $I$. Models serving this kind of role are often referred to as verifiers and are increasingly common in the literature on generative modeling [cobbe2021training, lightman2023let].
Our verifier is based on the Neural matrix factorization model [he2017neural]. Specifically, the verifier $v(U,I)$ approximates $\mathbb{P}(R{=}1\mid U,I)$ and is given by
$v(U,I)=g_{\gamma}\big(U,\texttt{CLIP}(I)\big),$ (7)
where $g_{\gamma}$ is a neural network, each user index $U$ is mapped to a trainable embedding vector, and $\texttt{CLIP}(\cdot)$ represents a fixed image embedding model [^2] We use OpenCLIP-ViT-bigG-14 [cherti2023reproducible], a different CLIP backbone than in Section Sec. 5.1, to avoid representation leak.. The verifier is trained by minimizing the binary cross-entropy loss [^3] This configuration achieves an ROC-AUC of approximately $87\%$ on both the training and test sets, with roughly $30\%$ of images in class 1. using the same training data employed to train the baselines and REBECA [^4] We use the training set to fit the verifier because it provides a larger number of labeled examples, allowing for a more accurate estimation of user preferences. While this could, in principle, introduce bias if both the verifier overfits, we find this unlikely in our case. As shown in Appendix E, the difference in the verifier’s performance on the training and test sets is not statistically different.. We then use the trained verifier $\hat{v}$ to quantify the personalization quality of generative models. Specifically, we define the personalization score of a generative model $\hat{p}(I\mid U,R=1)$ for user $U$ as
$\text{Score}\!\left(\hat{p}(I\mid U,R{=}1)\right)=\mathbb{E}_{\hat{p}(I\mid U,R{=}1)}[\,\hat{v}(U,I)\,],$ (8)
which we estimate empirically by sampling generated images for user $U$. This allows us to quantitatively compare different variants of REBECA and measure their ability to generate user-personalized images.
Aesthetics and Overall Quality. Personalization methods should preserve key aspects of image quality, including aesthetics and overall visual appeal. To evaluate whether REBECA and the baselines maintain high-quality generation while achieving personalization, we employ the LAION aesthetic predictor [laion_aesthetic_predictor] and the Human Preference Score (HPSv2) [wu2023human] to assess image aesthetics and perceptual quality. The two methods produce scalar scores that can be used to compare the quality of generated images.
5.4 Results
Original $t=0.1$ $t=0.2$ $t=0.3$ $t=0.4$
User 2
Refer to captionDisliked0.160.16 Refer to caption0.560.56 Refer to caption0.700.70 Refer to caption0.660.66 Refer to caption0.750.75
User 21
Refer to captionDisliked0.350.35 Refer to caption0.510.51 Refer to caption0.570.57 Refer to caption0.660.66 Refer to caption0.740.74
Figure 7: Interpolation results between latent representations (rows: users 2, 21). Leftmost image in each row was originally disliked (red badge). Columns show increasing $t\!\in\!\{0.1,0.2,0.3,0.4\}$ from the original. Overlay badges show the predicted preference score.
Precision@$k$ and Recall@$k$. Table 2 reports macro-averaged precision, recall, and F1 for 210 users. Across both @1 and @5, REBECA matches or surpasses all baselines. At lower prior CFG values the model attains the highest recall, sampling broadly from each user’s preference manifold but at lower precision. As CFG increases, generation concentrates around dominant preferences, raising precision while slightly reducing recall. Macro-F1 remains highest for REBECA, confirming the strongest overall precision-recall balance.
@1 @5
Model P R F1 P R F1
Shared LoRA 0.455 0.718 0.540 0.471 0.991 0.623
Per-user LoRA 0.485 0.713 0.557 0.470 0.986 0.621
VLM Persona 0.507 0.685 0.556 0.472 0.983 0.621
VLM Persona + LoRA 0.511 0.673 0.555 0.475 0.986 0.626
REBECA (CFG=$3.0$) 0.524 0.767 0.605 0.478 0.997 0.630
REBECA (CFG=$5.0$) 0.556 0.714 0.605 0.478 0.989 0.629
REBECA (CFG=$7.0$) 0.579 0.647 0.584 0.480 0.976 0.628
REBECA (CFG=$9.0$) 0.605 0.610 0.575 0.484 0.972 0.631
Table 2: Macro-averaged precision/recall/F1 over 210 users for retrieval of ground-truth likes from a liked+disliked pool using cosine similarity. Bold indicates best, underline indicates second-best per column.
Personalization. We compare its performance against several baseline methods. Figure 5 presents box-plots showing the distribution of user scores for each generation approach. REBECA clearly outperforms all alternatives, with VLM-based methods serving as strong baselines. We also observe that training a separate LoRA per user improves personalization compared to using a shared LoRA across users, though the gain is modest. This suggests that while LoRA introduces efficient low-rank adaptation, it lacks the inductive biases embedded in REBECA that are crucial for personalized image generation. In Appendix G, we further show disaggregated results.
Aesthetics and Overall Quality. Figure 6 reports results for both the LAION aesthetic predictor (left) and the HPSv2 human preference metric (right). The relative performance across methods closely mirrors what we observed in Figure 5, with REBECA consistently achieving the highest scores. Note that the HPSv2 metric is prompt-dependent: it evaluates the alignment between textual input and generated images. In our main evaluation (Fig. 6), we use the generic prompt to capture overall quality “Realistic image, finely detailed, with balanced composition and harmonious elements.” In Appendix H, we further evaluate with an empty prompt. In that setting, all generative methods exhibit similar overall quality, indicating that even in the absence of explicit textual guidance, REBECA maintains high image quality while delivering effective personalization.
[…]
Ablations. During generation, inference-time parameters such as classifier-free guidance (CFG) and system prompts may, in principle, influence the final output. We therefore analyze REBECA along two axes.
(i) Dual guidance. A joint sweep over the diffusion-prior CFG and the Stable Diffusion CFG (Fig. 8) reveals a clear trade-off between personalization strength and image fidelity. Increasing the prior CFG amplifies user-specific signal, while higher SD CFG improves aesthetic quality but mildly attenuates personalization. This interaction motivates our choice of moderate CFG values in the main experiments.
(ii) Prompt control. We further test whether prompt engineering can enhance output quality by evaluating three increasingly structured system prompts (see Sec. I.1). As shown in Fig. 9, all prompt levels yield nearly identical verifier and LAION scores across CFG settings, indicating that REBECA’s personalization is effectively prompt-independent. Even strong descriptive and negative prompts do not modulate the personalization signal, which instead resides in the latent user-conditioned embedding distribution learned by the diffusion prior, not in the text-conditioning pathway.
Figure 9: Prompt level for (a) verifier score and (b) LAION aesthetic quality. With REBECA, we do not observe a significant difference in scores with changing prompts
5.5 REBECA generates personalized images
We test whether REBECA’s gains are due solely to aesthetic quality by measuring personalization through a user-image matching experiment (Fig. 10). First, for each user, we compare the median verifier score of their own images to scores when images are randomly reassigned. The difference (correct minus random) quantifies personalized alignment. A permutation test [^5] Check the used algorithm in Appendix F. [lehmann1986testing] of $H_{0}:U\perp\!\!\!\!\perp I$ shows that correct matches score nearly five standard deviations above random (p $10^{-3}$, $\alpha=0.05$), confirming strong user-specific preference capture. Second, correlating verifier scores with average LAION aesthetic scores reveals a significant but modest relationship ($R^{2}=0.32$), indicating that aesthetic quality contributes partially to satisfaction but does not fully explain REBECA’s personalization.
Refer to caption Refer to caption
Figure 10: (a) Permutation test: correctly matched user-image pairs score nearly five standard deviations above random, indicating genuine personalization beyond aesthetics. (b) Aesthetic quality vs. verifier score.
6 Controllable Preference Navigation
REBECA embeddings enable controllable preference navigation via spherical interpolation between an original image and a personalized target. Given an image $I_{0}$ with embedding $z_{0}=\mathrm{CLIP}(I_{0})$ and a REBECA sample $z_{1}$ for a target user, we generate a trajectory using standard slerp [showmake_slerp, ramesh2022hierarchicaltextconditionalimagegeneration, levi2025the]:
$\mathrm{slerp}(z_{0},z_{1};t)=\frac{\sin((1-t)\theta)}{\sin\theta}\,\hat{z}_{0}+\frac{\sin(t\theta)}{\sin\theta}\,\hat{z}_{1},\quad t\in[0,1],$
where $\theta$ is the angle between normalized embeddings $\hat{z}_{0},\hat{z}_{1}$. Decoding these interpolated embeddings yields smooth edits that consistently increase the verifier’s predicted preference (Fig. 7), providing a user-aligned control axis orthogonal to text prompts.
7 Discussion
Discussion. REBECA demonstrates that implicit feedback alone can drive scalable personalization for diffusion models. By learning a lightweight conditional prior over CLIP embeddings and decoding with a frozen backbone, it provides a simple and general framework for connecting recommender signals with generative models—without paired comparisons, per-user fine-tuning, or LLM mediation. Because the method operates in embedding space, it naturally extends to other modalities such as audio, video, or cross-modal generators equipped with pretrained encoders.
Limitations include verifier shift arising from the surrogate preference model, the modest scale of our behavioral dataset, and cold-start scenarios where new users have limited interaction history. Future work includes reinforcement-style preference optimization directly on generative trajectories, large-scale implicit-feedback corpora, and richer multimodal conditioning to extend personalization beyond images.
Appendix A REBECA Prior Architecture
Our conditional diffusion prior is a lightweight and designed for personalized CLIP embeddings. Given a noisy embedding $I^{e}_{t}$ and conditioning, the model predicts the clean embedding $I^{e}_{0}$.
We first tokenize the CLIP embedding using a learned tokenizer that maps the 1D embedding into a small set of tokens. Each token is projected to a shared hidden dimension. Conditioning is injected through three tokens: user, rating, and timestep embeddings. These are mapped into a conditioning vector via a two-layer MLP.
The core of the model consists of $L$ PriorBlocks, each combining:
  • AdaLN-Zero layers for stable, scale-free conditioning,
  • self-attention over the image embedding tokens,
  • a gated cross-attention mechanism that selectively attends to the conditioning tokens, and
  • a zero-initialized MLP residual block for controlled feature updates.
All residual pathways are initialized at zero, ensuring training stability and preventing early over-conditioning. After $L$ blocks, the model projects tokens back to the original token dimension and merges them to reconstruct the predicted embedding. The entire architecture contains only 4.4M parameters, so each full training run takes under 10 minutes on a single RTX 4090.
Appendix B Training Protocol
Appendix C Image Tagging Pipeline
We employ the open-source checkpoint llava-hf/llava-1.5-7b-hf loaded with transformers. Each image is processed with the following fixed instruction:
You are a tagging engine. Return ONLY a JSON object with fields:
{"caption": str, "objects": [str], "attributes": [str],
 "styles": [str], "colors": [str]}.
Rules: lowercase; <=10 items total; no nulls.
We use the chat template provided by LLaVA for a single user turn …
Appendix D Baseline Specification
D.1 LoRA per User
Table 3: Per-user LoRA training hyperparameters. The LoRA rank ($r$), scaling factor ($\alpha$), dropout, training steps, and warmup schedule are adapted based on the number of available user images ($N_{i}$).
$N_{i}$ range LoRA $r$ LoRA $\alpha$ Dropout Steps Warmup
$N_{i}<8$ 8 8 0.10 1200 100
$8\leq N_{i}\leq 24$ 16 16 0.07 1500 120
$25\leq N_{i}\leq 60$ 32 32 0.05 1700 150
$61\leq N_{i}\leq 120$ 64 64 0.05 2000 200
$N_{i}>120$ 128 128 0.05 2500 200
D.2 Shared LoRA
D.3 VLM-Persona Generation
Prompt Construction.
Each user’s persona defines positive and negative keyword lists. When few positive terms exist, fallback categories (“nature, landscape, portrait, cityscape, animals”) are used. The builder cycles deterministically through available terms and appends either the user’s persona or a default stylistic tail.
Implementation.
We employ the Diffusers library for inference in half precision (fp16) on GPU, with the safety checker disabled for consistent reproducibility.
Table 4: Fallback prompt components used when user profiles lack sufficient detail.
Type Default values or description
Positive keywords {nature, landscape, portrait, cityscape, animals}
Style suffix “high quality, detailed, natural lighting”
Negative prompt {low quality, blurry, deformed, overexposed, underexposed}
Algorithm 1 Prompt Builder for User $u_{i}$
1:Input: profile $(\text{persona},\text{pos},\text{neg})$, image index $j$
[…]
4:if $P$ is empty then
5:  $P\leftarrow$ random fallback from Table 4
6:end if
[…]
10:return $(\text{prompt},\text{negprompt})$
Generation Summary.
  • Model: Stable Diffusion v1.5 (stable-diffusion-v1-5)
  • Inference steps: 50
  • Guidance scale: 5.0
  • Images per user: 25
  • Total users: 210
  • Seeding: $\text{seed}(u_{i},j)=42+10{,}000\times i+j$
  • Output: Serialized image bundles per user (.imgs)
This procedure, implemented in vlm_personas.ipynb, serves as the text-conditioned personalization baseline for comparison against LoRA-based and diffusion-prior models.
Appendix E Verifier diagnostics
Figure 11 shows that the verifier achieves nearly identical performance on the training and test sets. In the second panel, we report the results of a bootstrap analysis of the test data, demonstrating that the training ROC-AUC lies within two standard errors of the test ROC-AUC. This indicates that the verifier does not overfit and generalizes well to unseen samples.
Figure 11: Verifier performance on the training and test sets. The two curves are nearly indistinguishable, and the ROC-AUC values are not statistically different.
Appendix F Formally testing for personalization
Algorithm 2 implements a permutation test used to formally verify the personalization aspect of REBECA. In Algorithm 2, $\widehat{\text{Score}}$ is the median verifier score across users and $\widehat{\text{Score}_{b}}$ is the median verifier score across users after permutation using the random seed equals $b$.
Algorithm 2 Testing for REBECA personalization
1:Users, REBECA model, verifier $\hat{v}$, significance level $\alpha$, number of permutations $B=1000$
2:for each user do
3:  Generate 30 images using REBECA
4:end for
5:Compute baseline performance $\widehat{\text{Score}}$ using verifier $\hat{v}$
6:for $b=1$ to $B$ do
7:  Randomly permute generated images across users
8:  Compute $\widehat{\text{Score}}_{b}$ using verifier $\hat{v}$
9:end for
10:Compute p-value:
$p=\frac{1+\sum_{b=1}^{B}\mathds{1}\left[\widehat{\text{Score}}\leq\widehat{\text{Score}}_{b}\right]}{B+1}$
11:if $p\leq\alpha$ then
12:  Reject null hypothesis $H_{0}:U\perp\!\!\!\!\perp I$
13:  (that REBECA’s images do not depend on users)
14:end if
Appendix G Extra plots for the personalization results
Figures 12 and 13 present the performance of each method across users in terms of verifier scores and relative rankings, respectively. In both cases, REBECA consistently outperforms all baseline methods for the vast majority of users.
Figure 12: Verifier scores for each user across different generation methods. REBECA achieves the highest scores for most users, indicating stronger personalization performance.
Figure 13: Relative ranking of generation methods per user. REBECA ranks highest for the majority of users, outperforming all baselines in personalized generation quality.
Appendix H Extra plots for the aesthetics and overall quality results
Figure 14 reports the HPSv2 results obtained when using an empty prompt. In this zero-prompt setting, all generative methods display comparable overall quality. Notably, REBECA is capable of delivering personalized generations without hurting overall generation quality.
Figure 14: HPSv2 results under the zero-prompt setting. When no textual prompt is provided, all models achieve similar overall quality scores, indicating that prompt content mainly influences alignment rather than visual fidelity. REBECA maintains competitive quality while preserving personalization robustness.
Appendix I Ablations
I.1 System Prompts
To support the prompt–control ablation reported in the main paper, we evaluate REBECA under three system–prompt levels $c_{t}$, ranging from no prompt to strongly opinionated aesthetic prompts with aggressive negative prompts. For all experiments, we freeze the best-performing diffusion-prior checkpoint and generate personalized embeddings for all 210 users, which are then decoded using a Stable Diffusion v1.5 pipeline with an IP-Adapter (identical sampling settings across conditions).
The three prompt levels used in the ablation are:
Level 0: No prompt.
  • Positive prompt: ""
  • Negative prompt: ""
Level 1: Mild quality prompt.
  • Positive prompt: "high quality photo"
  • Negative prompt: "bad quality photo, letters"
Level 2: Strong descriptive/negative prompts.
  • Positive prompt: "Realistic image, finely detailed, with balanced composition and harmonious elements. Dynamic yet subtle tones, versatile style adaptable to diverse themes and aesthetics, prioritizing clarity and authenticity."
  • Negative prompt: "deformed, ugly, wrong proportion, frame, watermark, low res, bad anatomy, worst quality, low quality"
For each prompt level, we generate 10 personalized images per user (2,100 images per condition). The outputs are subsequently evaluated by our verifier model. As discussed in the main text, we observe no statistically significant effect of prompt level $c_{t}$ at any REBECA CFG value (Fig. 9), indicating that personalization arises from the diffusion prior rather than from prompt engineering.